# Polariton lattices as binarized neuromorphic networks

### v1

@@ Abstract
We introduce a novel neuromorphic network architecture based on a lattice of exciton-polariton condensates, intricately interconnected and energized through non-resonant optical pumping. The network employs a binary framework, where each neuron, facilitated by the spatial coherence of pairwise coupled condensates, performs binary operations. This coherence, emerging from the ballistic propagation of polaritons, ensures efficient, network-wide communication. The binary neuron switching mechanism, driven by the nonlinear repulsion through the excitonic component of polaritons, offers computational efficiency and scalability advantages over continuous weight neural networks. Our network enables parallel processing, enhancing computational speed compared to sequential or pulse-coded binary systems. The system’s performance was evaluated using the MNIST dataset for handwritten digit recognition, showcasing the potential to outperform existing polaritonic neuromorphic systems, as demonstrated by its impressive predicted classification accuracy of up to 97.5%.
@@ Introduction
The rapid development of artificial neural networks and applied artificial intelligence is predominantly aimed at the efficient processing of large data sets and pattern recognition. Traditional approaches, however, are increasingly facing constraints in terms of computational speed and energy efficiency, particularly in hardware implementations of these networks [1]. These constraints have spurred interest in neuromorphic systems, where hardware mimics the structure and function of the human brain. The exploration of novel materials and mechanisms is crucial for the development of efficient neuromorphic systems [2].
@@
Reflecting the latter statement, the authors of Ref. [3] leveraged the polariton properties to achieve a 93% level in recognizing handwritten digits in the Mixed National Institute of Standards and Technology dataset (MNIST dataset), that is a benchmark in pattern recognition tasks. Complementing this, Ref. [4] reported not only a 96.2% classification accuracy on the same dataset in the experiment but also demonstrated the efficiency of backpropagation training in their exciton-polariton-based neuromorphic hardware. This efficiency in processing complex patterns, however, often comes with increased computational and memory demands in traditional neuromorphic networks [11], which necessitates either a reduction in input resolution or more complex processing architectures for feasible operation times.
@@
In this manuscript, we present a neuromorphic network architecture employing lattices of exciton-polariton condensates, interconnected and energized through non-resonant optical pumping. This design capitalizes on benefits of a binary framework, wherein each neuron, aided by the spatial coherence of coupled condensates, executes binary operations. The lattice structure facilitates parallel uploading and processing of signals, enhancing neuron-neuron interactions. The effectiveness of our system was evaluated using the MNIST dataset. It demonstrated promising results compared to current neuromorphic systems. Additionally, we have developed a technique for input signal densing that notably improves the system performance, that achieves an accuracy rate of 97.5%, surpassing that of existing polariton-based neuromorphic systems.
@@
Figure 1: (a) A diagram depicting the possible experimental implementation of a lattice of pairwise coupled polariton condensates with optically controlled connections. Red cones symbolize input signal beams, and red arrows indicate the connections influenced by these beams. (b) A streamlined illustration of the envisaged structure, with gray, red, and blue circles denoting condensate lattice nodes, input, and output optical signals, respectively. Empty (filled) circles correspond to the absence (presence) of the signals. (c–e) Illustration of polariton dyad excitation in a planar microcavity, showing (c) profiles of two non-resonant optical pump spots for dyad excitation, (d,e) the condensates in the dyad with even (OFF) and odd (ON) interference patterns. (f-l) Depiction of the excitation of two adjacent dyads in OFF (f,h,k) and ON (g,j,l) states. Each pair of panels shows polariton density distribution (left) and pump intensity profiles, including the potential barrier (right). Switching between OFF and ON states is achieved using a signal optical pump beam equidistant from the four nodes. Panels (f,g) illustrate no separation between the dyads, (h,j) and (k,l) show the dyads separated by real (orange rectangle) and imaginary (blue rectangle) potential barriers, respectively.
Figure 2: Conceptual diagram of a binarized neural network based on a lattice of pairwise coupled polariton condensates. The initial signal originates from a grayscale image from the MNIST dataset, which is binarized and projected onto a $n_{\text{in}}\times n_{\text{in}}$ transformation lattice. This lattice serves as a pattern for the input optical signal. This signal then activates neurons within the reservoir lattice, generating the resultant optical output signal. Subsequently, the output is processed via a linear classifier (LC).
@@ Results
As observed in the geometry presented in Figs. 1(f,g), the condensates are interconnected not only within each dyad but also with other condensates belonging to an adjacent dyad. In such a configuration, it’s more appropriate to refer to a tetrad, involving all condensates excited within an elementary cell. Moreover, in the limit of a full-size lattice, due to the macroscopic spatial coherence of polariton condensates, such connections may extend well beyond a single cell, suggesting a complex network of interactions across the lattice structure. To maintain the paradigm of pairwise interactions, it is essential to isolate the dyads from each other within the polariton lattice. This can be achieved by introducing additional potential barriers between condensates whose connections need to be severed. Approaches to effectively isolating individual dyads will be discussed in the Discussion section. Figures 1(h) and 1(j) illustrate pairs of polariton dyads in OFF and ON states, respectively, with the dyads separated by real potential barriers. In Figs. 1(k) and 1(l), imaginary absorbing barriers are used for the separation. It is evident that independently of the barrier’s nature, the controlling optical beam, in both scenarios, switches both dyads from the OFF state to the ON state, while maintaining the dyads as non-interacting entities, ensuring that their interconnections remain unaffected.
Figure 3: Schematic depicting the conversion of a binarized initial signal lattice of size $n_{0}\times n_{0}$ into an input signal lattice of size $n_{\text{in}}\times n_{\text{in}}$, utilizing randomization and expansion, both without (a) and with (b) signal densing.
@@
Typically, a neuromorphic network comprises in its core an input layer where signals are initially received, processed through a series of intermediate hidden layers. The output layer then finalizes the processing sequence. In this paper, we propose a binarized polariton network with a single hidden layer based on a lattice of polariton condensates, see Fig. 1(a). For the sake of simplicity, we consider square lattices, characterized by an equal number of nodes $n_{\text{c}}$ along each side, although our results can be generalized to lattices of an arbitrary shape. Within the lattice, adjacent condensates are randomly paired into dyads, that are mutually isolated, ensuring that the state of each dyad remains independent and it does not affect adjacent pairs. The dyads in the network function as binary neurons. They switch between OFF and ON states, corresponding to the output signal values of 0 and 1, respectively, in response to the absence or presence of an additional pumping control signal near each dyad, which acts as an input signal. Each dyad at the center of the lattice can be subjected to two input signals, while dyads at the edges are influenced by only one signal, cf. (i) and (ii–iv) in Fig. 1(a). The dyad unaffected by the input signal (ii) remains in the OFF state, while the dyads affected by one (iii) or two (iv) input signals switch to the ON state. Due to the randomness in pairing condensates within the lattice, some individual condensates may not be part of any dyads, see (v) in Fig. 1(a). In such cases, these condensates do not contribute to the generation of the output signal.
@@
In our proposal, the inherent randomness in the pairing of adjacent condensates into dyads within the lattice introduces a significant element of nonlinearity into the network’s functioning. This randomness in dyad formation means that the impact of input signals on the neurons varies, with some dyads being activated by one or two signals, while others remain unaffected. This randomness leads to a complex network response that cannot be easily mapped or predicted linearly. Such variability in dyad responses, triggered by input signals, enhances the computational capabilities of the hidden layer. The dyads, functioning as binary neurons, switch between OFF and ON states in response to control signals, similar to a classical OR gate, where receiving at least one “1” input signal triggers a “1” output signal. However, the diverse arrangement of these dyads, some isolated and some interconnected, ensures a dynamic and non-linear processing environment within the network, crucial for tackling intricate computational tasks.
@@
On the next step, the pattern preparation takes place, which involves several operations simultaneously, see Fig. 3(a). The first operation is randomization. The elements of the binarized matrix are transferred to the pattern matrix not sequentially, but in a random manner, using a randomization mask that is consistent for all recognized images. This technique allows for a more uniform distribution of the signal across the entire layer, engaging both central and peripheral neurons to the same extent. Such an approach ensures that the neural network utilizes its entire structure more effectively, enhancing overall performance and accuracy.
@@
The second operation involves expansion, which entails increasing the number of neurons engaged in processing. This step enhances the network’s capacity to handle and interpret complex data by involving a larger array of neurons, thereby improving its computational power and efficiency. In the geometry of the square lattice under consideration, the number of neurons in the hidden layer should increase superlinearly with the number of elements in the input signal lattice, $N_{\text{in}}=n_{\text{in}}^{2}$, as $N_{\text{d}}\leq N_{\text{in}}+2\sqrt{N_{\text{in}}}+1$. To increase the lattice size from $n_{0}$ to $n_{\text{in}}$, elements randomly selected from the binarized lattice are repeatedly inserted into the input pattern lattice, see (iii) and (iv) in Fig. 3. One should note that both randomization and expansion are uniformly applied to all initial images, ensuring that each element of the binarized matrix is allocated to specific, unchanging positions within the input pattern lattice, consistent across different images. This approach to element placement standardizes the processing framework for each image.
@@
Figure 4: Accuracy of the MNIST handwritten digit recognition in dependence on the size of the input signal lattice $n_{\text{in}}$ (lower scale) or the number of neurons (dyads) in the hidden layer $N_{\text{d}}$ (upper scale). Each data point represents an average from ten numerical experiments, each utilizing different randomization masks. The shaded area reflects the variation in accuracy across these numerical experiments. Vertical lines, serving as guides for the eye, indicate the conditions, where the size of the polariton lattice matches that of the initial images (dashed), and when the number of neurons equals the number of pixels in the initial image (dash-dotted).
@@
To evaluate the effectiveness of the proposed polariton neuromorphic network architecture, we simulated its operation using numerical calculations. The assessment was conducted on the MNIST dataset, comprising a set of 60000 samples of handwritten digits for training the network and a set of 10000 samples for testing purposes. Given that the MNIST dataset serves as a standard benchmark for the goals of image classification, its use in our study enables a rigorous comparison of our findings with existing results in polariton-based neuromorphic network research [3, 4, 12, 13]. It is noteworthy that the binary nature of both the input and output signals, as well as the functioning of the neurons within the hidden layer, facilitated a more efficient utilization of computational resources. This efficiency, in turn, permitted the deployment of the entire spectrum of training and testing images, eliminating the need for any artificial reduction in image resolution. At the linear classifier stage of the neural network, the linear regression algorithm was employed for efficient data classification.
@@
It is commonly observed in the neuromorphic network research that the increase of the number of neurons in the hidden layer can potentially improve the classification accuracy. This is corroborated by our findings, as illustrated in Fig. 4, where we demonstrate a monotonic increase in the accuracy of classifying MNIST dataset images with the increase in the size of the input signal lattice $n_{\text{in}}$, which implies an increase in the number of neurons $N_{\text{d}}$ in the hidden layer. In the figure, each data point corresponds to the accuracy of the image recognition at a given number of neurons, averaged over ten numerical experiments, carried out for different randomisation masks used at the input layer preparation step. The light blue shaded area indicates the range of variation of accuracy obtained in the corresponding numerical experiment series. It can be observed that randomization has a finite effect on accuracy, altering it within a margin of less than one percent. This dependence exhibits a saturating character, converging towards about 95.4% accuracy with an increasing number of neurons. Vertical lines indicate the observations, corresponding to the size of the input layer equal to the size of the initial image, ${n_{\text{in}}=n_{0}=28}$, with the number of neurons estimated as about 420 (gray dashed), and to the size of the input layer ${n_{\text{in}}=39}$, with the number of neurons close to the number of pixels in the initial image, $N_{\text{d}}=760<n_{0}^{2}$ (black dot-dashed). It can be seen that in the second case, even with a slightly smaller number of neurons involved, the classification accuracy of the polariton neuromorphic network, averaging over 93.3%, exceeds the accuracy of the software linear classification of both grayscale (92.5%) and binarized (91.9%) images. In addition, reducing the input signal lattice dimension to 20, and consequently the number of neurons in the hidden layer to approximately 220, still ensures an accuracy above 90%.
Figure 5: The classification accuracy in the dependence of the overlap degree $s$ of the input signal for square polariton lattice systems of different size $d_{\text{in}}$ with different numbers of neurons in the interaction layer: $d_{\text{in}}=20$ with 220 neurons (blue), $d_{\text{in}}=28$ with 420 neurons (green), $d_{\text{in}}=39$ with 760 neurons (red), $d_{\text{in}}=51$ with 1350 neurons (violet) and $d_{\text{in}}=160$ with 12960 neurons (brown). Red stars indicate the maxima of the dependencies. Dashed lines indicate the accuracy levels for alternative established classification approaches: the linear software classification of the grayscale MNIST data set (92.5%), the binarized data MBIST set (91.9%), the binarized polariton network based on a layer of XOR gates [12] and the nonlinear polariton network with the software backpropagation training [4].
@@
In the approach not involving the image densing, all elements can be filled consecutively. Once a particular element in the input layer is assigned a value, this value remains unchanged, and the element no longer participates in the further process of filling the input layer. Each such element contains only one incoming connection from the projected binarized layer, see blue lines in Fig. 3(a). In contrast, in the case of image densing, the value of a filled element can be modified during the subsequent process. This approach is illustrated in Fig. 3(b). If two elements of the binarized image matrix are used for determining an element of the input matrix, in this case, the principle of assigning values is similar to the operation of a logical OR gate. If both elements are 0, the assigned element in the input layer also takes a value 0, see (v) in Fig. 3(b). If either one of the elements (vi) or both of them (vii) are 1, then the assigned element is also 1. This can be easily generalized to the case of three and more elements.

### v2

@@ Abstract
We introduce a novel neuromorphic network architecture based on a lattice of exciton-polariton condensates, intricately interconnected and energized through non-resonant optical pumping. The network employs a binary framework, where each neuron, facilitated by the spatial coherence of pairwise coupled condensates, performs binary operations. This coherence, emerging from the ballistic propagation of polaritons, ensures efficient, network-wide communication. The binary neuron switching mechanism, driven by the nonlinear repulsion through the excitonic component of polaritons, offers computational efficiency and scalability advantages over continuous weight neural networks. Our network enables parallel processing, enhancing computational speed compared to sequential or pulse-coded binary systems. The system’s performance was evaluated using diverse datasets, including the MNIST dataset for image recognition and the Speech Commands dataset for voice recognition tasks. In both scenarios, the proposed system demonstrates the potential to outperform existing polaritonic neuromorphic systems. For image recognition, this is evidenced by an impressive predicted classification accuracy of up to 97.5%. In voice recognition, the system achieved a classification accuracy of about 68% for the ten-class subset, surpassing the performance of conventional benchmark, the Hidden Markov Model with Gaussian Mixture Model.
@@ Introduction
The rapid development of artificial neural networks and applied artificial intelligence is predominantly aimed at the efficient processing of large data sets and pattern recognition [1]. Traditional approaches, however, are increasingly facing constraints in terms of computational speed and energy efficiency, particularly in hardware implementations of these networks [2]. These constraints have spurred interest in neuromorphic systems, where hardware mimics the structure and function of the human brain. The exploration of novel materials and mechanisms is crucial for the development of efficient neuromorphic systems [3].
@@
Reflecting the latter statement, the authors of Ref. [4] leveraged the polariton properties to achieve a 93% level in recognizing handwritten digits in the Mixed National Institute of Standards and Technology dataset (MNIST dataset) [12, 13], that is a benchmark in pattern recognition tasks. Complementing this, Ref. [5] reported not only a 96.2% classification accuracy on the same dataset in the experiment but also demonstrated the efficiency of backpropagation training in their exciton-polariton-based neuromorphic hardware. This efficiency in processing complex patterns, however, often comes with increased computational and memory demands in traditional neuromorphic networks [14], which necessitates either a reduction in input resolution or more complex processing architectures for feasible operation times.
@@
In this manuscript, we present a neuromorphic network architecture employing lattices of exciton-polariton condensates, interconnected and energized through non-resonant optical pumping. This design capitalizes on benefits of a binary framework, wherein each neuron, aided by the spatial coherence of coupled condensates, executes binary operations. The lattice structure facilitates parallel uploading and processing of signals, enhancing neuron-neuron interactions. The effectiveness of our system was evaluated using the MNIST dataset. It demonstrated promising results compared to current neuromorphic systems. Additionally, we have developed a technique for input signal densing that notably improves the system performance, that achieves an accuracy rate of 97.5%, surpassing that of existing polariton-based neuromorphic systems. To further validate our model’s robustness and versatility, we extended our testing to the Speech Commands dataset [37, 38], which includes audio clips for voice command recognition. Our system’s performance, considerably exceeding that of both the linear classifier and the advanced Hidden Markov Model with Gaussian Mixture Model (HMM-GMM) [39, 40], confirms its potential for effective application in complex speech recognition tasks.
@@
Figure 1: (a) A diagram depicting the possible experimental implementation of a lattice of pairwise coupled polariton condensates with optically controlled connections. Red cones symbolize input signal beams, and red arrows indicate the connections influenced by these beams. The dyads are numbered with Roman numerals to correspond with the numbering in Fig. 2. (b) A streamlined illustration of the envisaged structure, with gray, red, and blue circles denoting condensate lattice nodes, input, and output optical signals, respectively. Empty (filled) circles correspond to the absence (presence) of the signals. (c–e) Illustration of polariton dyad excitation in a planar microcavity, showing (c) profiles of two non-resonant optical pump spots for dyad excitation, (d,e) the condensates in the dyad with even (OFF) and odd (ON) interference patterns. (f-l) Depiction of the excitation of two adjacent dyads in OFF (f,h,k) and ON (g,j,l) states. Each pair of panels shows polariton density distribution (left) and pump intensity profiles, including the potential barrier (right). Switching between OFF and ON states is achieved using a signal optical pump beam equidistant from the four nodes. Panels (f,g) illustrate no separation between the dyads, (h,j) and (k,l) show the dyads separated by real (orange rectangle) and imaginary (blue rectangle) potential barriers, respectively.
Figure 2: Depiction of excitation of a lattice of dyads of geometry, schematically shown in Figs. 1(a) and 1(b), in the absence (a) and in the presence (b) of control signals. The panels show time-integrated spatial distribution of the polariton density. Regions with densities higher than the range covered by the color scale are indicated in white. (c) The profiles of the non-resonant pump spots (main color scheme), the trapping potential profile for isolating dyads (white) and the control beam profiles (blue) for toggling dyads between ON and OFF states. The indices (i) to (xii) enumerate neurons in (a) and (b). (d) The dependence of distinguishability, $\Delta\tilde{I}$, of the OFF and ON signals on the intensities of the pump pulses for condensates in dyads, $P_{10}$, and intensities of signal pulses, $P_{20}$. Values of $\Delta\tilde{I}<0$ are colored in black. Pulse durations are taken as $w_{\tau 1}=5$ ps and $w_{\tau 2}=8$ ps, respectively. Definition of the distinguishability, $\Delta\tilde{I}$, is given in the text. Star markers enumerated from 1 to 17 indicate maximal $\Delta\tilde{I}$ at given $P_{10}$. The red star indicates the parameters used for (a) and (b). (e) Variation of the relative intensity of output signals, $\bar{I}$, in each neuron from (i) to (xii) at the pump intensities $(P_{10},P_{20})$ corresponding to points $j=1,2,...17$ in (d). In (iv), the dependence should be multiplied by 2.5. Each dot corresponds to a separate numerical experiment, in which positions of the pump pulses, that excite condensates in dyads across the lattice, deviate randomly from their owing positions in the range of distances from $-\delta$ to $+\delta$. Dots of different colors correspond to different deviations $\delta$. Gray lines used as references indicate the average of the minimal relative intensity of the neuron in the ON state and the maximal relative intensity of the neuron in the OFF state in the absence of deviations of the pump pulses.
@@ Results
Figure 3: Conceptual diagram of a binarized neural network based on a lattice of pairwise coupled polariton condensates. The initial signal originates from a grayscale image from the MNIST dataset, which is binarized and projected onto a $n_{\text{in}}\times n_{\text{in}}$ transformation lattice. This lattice serves as a pattern for the input optical signal. This signal then activates neurons within the hidden layer, generating the resultant optical output signal. Subsequently, the output is processed via a linear classifier (LC).
@@
As observed in the geometry presented in Figs. 1(f,g), the condensates are interconnected not only within each dyad but also with other condensates belonging to an adjacent dyad. In such a configuration, it’s more appropriate to refer to a tetrad, involving all condensates excited within an elementary cell. Moreover, in the limit of a full-size lattice, due to the macroscopic spatial coherence of polariton condensates, such connections may extend well beyond a single cell, suggesting a complex network of interactions across the lattice structure. To maintain the paradigm of pairwise interactions, it is essential to isolate the dyads from each other within the polariton lattice. This can be achieved by introducing additional potential barriers between condensates whose connections need to be severed. Figures 1(h) and 1(j) illustrate pairs of polariton dyads in OFF and ON states, respectively, with the dyads separated by real potential barriers. In Figs. 1(k) and 1(l), imaginary absorbing barriers are used for the separation. It is evident that independently of the barrier’s nature, the controlling optical beam, in both scenarios, switches both dyads from the OFF state to the ON state, while maintaining the dyads as non-interacting entities, ensuring that their interconnections remain unaffected.
While managing isolation for a single pair of dyads is relatively straightforward, achieving the same for a full lattice of dyads requires a more comprehensive approach. This involves carefully selecting the appropriate nature and configuration of the potentials to provide sufficient isolation for each dyad on all sides without impeding their interaction with control signals. Figure 2 presents the results of numerical simulations for a lattice of dyads, schematically depicted in Figs. 1(a) and 1(b), both in the absence [Fig. 2(a)] and in the presence [Fig. 2(b)] of signal pulses. The shape of the potential that ensures the successful operation of each artificial neuron within the lattice is schematically depicted in white in Fig. 2(c). A more detailed analysis of the resulting behavior of artificial neurons under various isolation conditions and interactions with control pulses is provided in Sec. S3 of the Supplementary Information. The possible experimental implementation of the proposed system is discussed in Sec. S6 of the Supplementary Information.
@@
Typically, a neuromorphic network comprises in its core an input layer where signals are initially received, processed through a series of intermediate hidden layers. The output layer then finalizes the processing sequence. In this paper, we propose a binarized polariton network with a single hidden layer based on a lattice of polariton condensates, see Fig. 1(a). For the sake of simplicity, we consider square lattices, characterized by an equal number of nodes $n_{\text{c}}$ along each side, although our results can be generalized to lattices of an arbitrary shape. Within the lattice, adjacent condensates are randomly paired into dyads, that are mutually isolated, ensuring that the state of each dyad remains independent and it does not affect adjacent pairs. The dyads in the network function as binary neurons. They switch between OFF and ON states, corresponding to the output signal values of 0 and 1, respectively, in response to the absence or presence of an additional pumping control signal near each dyad, which acts as an input signal. Each dyad at the center of the lattice can be subjected to two input signals, while dyads at the edges are influenced by only one signal, cf. (x) and other numbered dyads in Fig. 1(a). The dyad unaffected by the input signal (vii) remains in the OFF state, while the dyads affected by one (v) or two (iv) input signals switch to the ON state. Due to the randomness in pairing condensates within the lattice, some individual condensates may not be part of any dyads, see (xiii) in Fig. 1(a). In such cases, these condensates do not contribute to the generation of the output signal.
@@
In our proposal, the inherent randomness in the pairing of adjacent condensates into dyads within the lattice introduces a significant element of nonlinearity into the network’s functioning. This randomness in dyad formation means that the impact of input signals on the neurons varies, with some dyads being activated by one or two signals, while others remain unaffected. This randomness leads to a complex network response that cannot be easily mapped or predicted linearly. Such variability in dyad responses, triggered by input signals, enhances the computational capabilities of the hidden layer. The dyads, functioning as binary neurons, switch between OFF and ON states in response to control signals, similar to a classical OR gate, where receiving at least one ‘1’ input signal triggers a ‘1’ output signal. However, the diverse arrangement of these dyads, some isolated and some interconnected, ensures a dynamic and non-linear processing environment within the network, crucial for tackling intricate computational tasks. For a more detailed discussion of the nonlinearity induced by the structure of the neural network, see Sec. S1 of the Supplemental Information.
Using the hidden layer configuration illustrated in Fig. 1(a,b) as an example, we demonstrate the feasibility of practical implementation of a hidden layer based on a lattice of polariton condensates. Using a model based on the generalized Gross-Pitaevskii equation for the polariton wave function $\Psi(t,\mathbf{r})$, as detailed in the Methods section, we conduct a series of numerical experiments to simulate the excitation of a lattice of polariton dyads under control by laser pulses. The polariton condensates are excited by non-resonant laser pulses with intensity ${P_{1}(t,\mathbf{r})=P_{10}f_{1}(t,\mathbf{r},w_{1},w_{\tau 1})}$, arranged in a $5\times 5$ lattice configuration as shown in Fig. 1(a,b). Here, $P_{10}$ is the intensity magnitude, and $f_{1}(t,\mathbf{r},w_{1},w_{\tau 1})$ is a Gaussian function of space and time with parameters of width $w_{1}$ and duration $w_{\tau 1}$ of pulses. The signal pulses with intensity ${P_{2}(t,\mathbf{r})=P_{20}f_{2}(t,\mathbf{r},w_{2},w_{\tau 2})}$ are applied to specific nodes in the intervals between the condensate pump pulses. The meaning of the parameters in ${P_{2}(t,\mathbf{r})}$ is analogous to that in ${P_{1}(t,\mathbf{r})}$. The positions of the pump spots are indicated in Fig. 2(c). Shapes of pulses as well as values of the parameters used for simulations are detailed in the Methods section.
For the mutual isolation of dyads, we employ an effective potential, schematically depicted in white in Fig. 2(c). It is important to note, however, that despite its sophisticated shape, there is no need for a unique potential design for each dyad. The shape of the potential within each dyad is identical for all dyads, and the final lattice configuration is assembled from these uniform building blocks. Approaches to forming localization potentials for polaritons are discussed in the Discussion section.
[…]
Since the pump intensity gradually increases in the pulsed excitation regime, dyads switch to the ON state gradually rather than instantaneously, and they may change their parity during evolution. Consequently, even in the OFF state, the time-integrated intensity at the center of the dyad can be non-zero. However, it must always remain lower than the intensity of a dyad in the ON state. To analyze whether the pump parameters meet this condition, it is useful to introduce the parameter of distinguishability of OFF and ON states, $\Delta\tilde{I}$. For a polariton lattice with a given configuration, $\Delta\tilde{I}$ represents the difference between the minimum intensity of output signals from dyads expected to be in the ON state, $\text{Min}(I_{\text{ON}})$, and the maximum intensity of signals from dyads expected to be in the OFF state, $\text{Max}(I_{\text{OFF}})$, normalized by the latter, $\Delta\tilde{I}=[\text{Min}(I_{\text{ON}})-\text{Max}(I_{\text{OFF}})]/\text{Max}(I_{\text{OFF}})$. A successful choice of pump parameters is characterized by positive values of $\Delta\tilde{I}$. If $\Delta\tilde{I}$ is zero or negative, the ON and OFF signals become indistinguishable. The color map in Fig. 2(d) outlines the region in the plane of pump pulse intensities $(P_{10},P_{20})$ where ON and OFF states of the artificial neurons are distinguishable, with other parameters held constant. It is evident that, under the optimal combination of parameters (indicated by the blue line in the panel), the distinguishability $\Delta\tilde{I}$ can reach a value of one, meaning that the difference in intensities is comparable to the intensity of the output signal of the neuron in the OFF state. Such distinguishability is expected to provide excellent robustness against noise. It significantly exceeds the corresponding measure in [15], which, according to rough estimates, does not exceed 0.2 for a single neuron.
The distribution of polariton density within each dyad represents a time-averaged interference pattern, largely determined by the relative positions of the condensates within the dyad and the localization potential. We investigated the sensitivity of the output signal of the polariton lattice to deviations in the positions of the condensates from their designated positions. For this purpose, for several pairs of pump pulse intensities, marked in Fig. 2(d) (numbered from $j=1$ to $j=17$), we calculated the output signal intensities of the neurons in the treated example lattice configuration, in the presence of input signals. In Fig. 2(e), we present the normalized output signal intensities for all neurons as a function of $j$. The normalization is done with respect to the reference intensity, which for each neuron is the output signal intensity in the absence of the input signal. For each $j$, a series of independent numerical experiments was conducted, where the positions of the pump spots were randomly deviated from their designated positions within the range from $-\delta$ to $+\delta$. Each point on the individual panels in Fig. 2(e) corresponds to a separate numerical experiment, with different colors representing different magnitudes of deviation $\delta$. It can be observed that random deviations within approximately ±$\pm 0.5\,\mu\text{m}$ do not undermine the distinguishability of the neuron states.
@@
On the next step, the pattern preparation takes place, which involves several operations simultaneously, see Fig. 4(a1). The first operation is randomization. The elements of the binarized matrix are transferred to the pattern matrix not sequentially, but in a random manner, using a randomization mask that is consistent for all recognized images. This technique allows for a more uniform distribution of the signal across the entire layer, engaging both central and peripheral neurons to the same extent. Such an approach ensures that the neural network utilizes its entire structure more effectively, enhancing overall performance and accuracy.
@@
The second operation involves expansion, which entails increasing the number of neurons engaged in processing. This step enhances the network’s capacity to handle and interpret complex data by involving a larger array of neurons, thereby improving its computational power and efficiency. In the geometry of the square lattice under consideration, the number of neurons in the hidden layer should increase superlinearly with the number of elements in the input signal lattice, $N_{\text{in}}=n_{\text{in}}^{2}$, as $N_{\text{d}}\leq(N_{\text{in}}+2\sqrt{N_{\text{in}}}+1)/2$. To increase the lattice size from $n_{0}$ to $n_{\text{in}}$, elements randomly selected from the binarized lattice are repeatedly inserted into the input pattern lattice, see (iii) and (iv) in Fig. 4(a1). One should note that both randomization and expansion are uniformly applied to all initial images, ensuring that each element of the binarized matrix is allocated to specific, unchanging positions within the input pattern lattice, consistent across different images. This approach to element placement standardizes the processing framework for each image.
@@
Figure 4: Evaluation of the MNIST handwritten digit recognition by the polariton neuromorphic network. (a1,a2) Schematic depicting the conversion of a binarized initial signal lattice of size $n_{0}\times n_{0}$ into an input signal lattice of size $n_{\text{in}}\times n_{\text{in}}$, utilizing randomization and expansion, both without (a1) and with (a2) signal densing. (b) The recognition accuracy in dependence on the size of the input signal lattice $n_{\text{in}}$ (lower scale) or the number of neurons (dyads) in the hidden layer $N_{\text{d}}$ (upper scale). Each data point represents an average from ten numerical experiments, each utilizing different randomization masks. The shaded area reflects the variation in accuracy across these numerical experiments. Vertical lines, serving as guides for the eye, indicate the conditions, where the size of the polariton lattice matches that of the initial images (dashed), and when the number of neurons equals the number of pixels in the initial image (dash-dotted). (c) The recognition accuracy in dependence of the densing degree $s$ of the input signal for square polariton lattice systems of different size $d_{\text{in}}$ with different numbers of neurons in the interaction layer: $d_{\text{in}}=20$ with 220 neurons (blue), $d_{\text{in}}=28$ with 420 neurons (green), $d_{\text{in}}=39$ with 800 neurons (red), $d_{\text{in}}=75$ with 2890 neurons (violet), and $d_{\text{in}}=160$ with 12960 neurons (brown). Red markers indicate the maxima of the dependencies. Horizontal dashed lines indicate the accuracy levels for alternative established classification approaches (from bottom to top): the linear software classification of the grayscale (92.5%) and binarized (91.9%) MNIST data set, the binarized polariton network based on a layer of XOR gates [15] and the nonlinear polariton network with the software backpropagation training [5].
@@
To evaluate the effectiveness of the proposed polariton neuromorphic network architecture, we simulated its operation using numerical calculations. The assessment was conducted on the MNIST dataset [12, 13], comprising a set of $60000$ samples of handwritten digits for training the network and a set of $10000$ samples for testing purposes. Given that the MNIST dataset serves as a standard benchmark for the goals of image classification, its use in our study enables a rigorous comparison of our findings with existing results in polariton-based neuromorphic network research [4, 5, 15, 16]. It is noteworthy that the binary nature of both the input and output signals, as well as the functioning of the neurons within the hidden layer, facilitated a more efficient utilization of computational resources. This efficiency, in turn, permitted the deployment of the entire spectrum of training and testing images, eliminating the need for any artificial reduction in image resolution. At the linear classifier stage of the neural network, the linear regression algorithm was employed for efficient data classification.
@@
It is commonly observed in the neuromorphic network research that the increase of the number of neurons in the hidden layer can potentially improve the classification accuracy. This is corroborated by our findings, as illustrated in Fig. 4(b), where we demonstrate a monotonic increase in the accuracy of classifying MNIST dataset images with the increase in the size of the input signal lattice $n_{\text{in}}$, which implies an increase in the number of neurons $N_{\text{d}}$ in the hidden layer. In the figure, each data point corresponds to the accuracy of the image recognition at a given number of neurons, averaged over ten numerical experiments, carried out for different randomisation masks used at the input layer preparation step. The light blue shaded area indicates the range of variation of accuracy obtained in the corresponding numerical experiment series. It can be observed that randomization has a finite effect on accuracy, altering it within a margin of less than one percent. This dependence exhibits a saturating character, converging towards about 95.4% accuracy with an increasing number of neurons. Vertical lines indicate the observations, corresponding to the size of the input layer equal to the size of the initial image, ${n_{\text{in}}=n_{0}=28}$, with the number of neurons estimated as about 420 (gray dashed), and to the size of the input layer ${n_{\text{in}}=39}$, with the number of neurons close to the number of pixels in the initial image, $N_{\text{d}}=760<n_{0}^{2}$ (black dot-dashed). It can be seen that in the second case, even with a slightly smaller number of neurons involved, the classification accuracy of the polariton neuromorphic network, averaging over 93.3%, exceeds the accuracy of the software linear classification of both grayscale (92.5%) and binarized (91.9%) images, see also Sec. S2 in the Supplementary Information.
In addition, reducing the input signal lattice dimension to 20, and consequently the number of neurons in the hidden layer to approximately 220, still ensures an accuracy above 90%.
@@
In the approach not involving the image densing, all elements can be filled consecutively. Once a particular element in the input layer is assigned a value, this value remains unchanged, and the element no longer participates in the further process of filling the input layer. Each such element contains only one incoming connection from the projected binarized layer, see blue lines in Fig. 4(a1). In contrast, in the case of image densing, the value of a filled element can be modified during the subsequent process. This approach is illustrated in Fig. 4(a2). If two elements of the binarized image matrix are used for determining an element of the input matrix, in this case, the principle of assigning values is similar to the operation of a logical OR gate. If both elements are 0, the assigned element in the input layer also takes a value 0, see (v) in Fig. 4(a2). If either one of the elements (vi) or both of them (vii) are 1, then the assigned element is also 1. This can be easily generalized to the case of three and more elements.